# LoRA Done RITE: Robust Invariant Transformation Equilibration for LoRA Optimization

### v1

@@ 3.1 Diagonal preconditioning is not enough for transformation invariance
In this section we show that diagonal preconditioning is not enough to achieve transformation invariance. Recall that the LoRA factors ${\bm{A}}\in\mathbb{R}^{m\times r}$ and ${\bm{B}}\in\mathbb{R}^{n\times r}$. When updating ${\bm{A}}$, most existing optimizers utilize the following update:
$\operatorname{vec}(\delta{\bm{A}})=-{\bm{P}}\operatorname{vec}(\nabla{\bm{A}}),$ (9)
@@
where $\operatorname{vec}(\cdot)$ lists the elements of matrix in a vector in column-major order and ${\bm{P}}\in\mathbb{R}^{mr\times mr}$ is a symmetric positive definite preconditioning matrix. Diagonal preconditioning methods like Adam and Adagrad assume ${\bm{P}}$ is a diagonal matrix, while matrix preconditioning methods such as Shampoo and CASPR use non-diagonal ${\bm{P}}$.
Consider the LoRA pairs $({\bm{A}}_{1},{\bm{B}}_{1})$ and $({\bm{A}}_{2},{\bm{B}}_{2})$ such that:
${\bm{A}}_{2}={\bm{A}}_{1}{\bm{R}},\quad{\bm{B}}_{2}={\bm{B}}_{1}{\bm{R}}^{-\top},$
@@
where ${\bm{R}}$ is an invertible matrix. We note that the above is a matrix generalization of example (6). We can now derive the gradient with respect to LoRA factors using (7):
$\nabla{\bm{A}}_{2}=\nabla{\bm{Z}}{\bm{B}}_{2}=\nabla{\bm{A}}_{1}{\bm{R}}^{-\top}.$
@@
Let us consider the case when $n=m=1$. For ${\bm{A}}\in\mathbb{R}^{m\times r}$, we can rewrite $\operatorname{vec}(\delta{\bm{A}})^{\top}=\delta{\bm{A}}$ and $\operatorname{vec}(\nabla{\bm{A}})^{\top}=\nabla{\bm{A}}$ in update (9):
$\delta{\bm{A}}=-\nabla{\bm{A}}{\bm{P}}^{\top}\implies\delta{\bm{A}}{\bm{B}}^{\top}=-\nabla{\bm{A}}{\bm{P}}^{\top}{\bm{B}}^{\top}.$
@@
In this case, if we have two equivalent LoRA pairs $({\bm{A}}_{1},{\bm{B}}_{1}),({\bm{A}}_{2},{\bm{B}}_{2})$ with their corresponding preconditioners ${\bm{P}}_{1},{\bm{P}}_{2}$, then
$\delta{\bm{A}}_{2}{\bm{B}}_{2}^{\top}=-\nabla{\bm{A}}_{2}{\bm{P}}_{2}^{\top}{\bm{B}}_{2}^{\top}=-\nabla{\bm{A}}_{1}({\bm{R}}^{-\top}{\bm{P}}_{2}{\bm{R}}^{-1}){\bm{B}}_{1}^{\top}.$ (10)
If transformation invariance holds, then:
$\delta{\bm{A}}_{1}{\bm{B}}_{1}^{\top}=-\nabla{\bm{A}}_{1}{\bm{P}}_{1}^{\top}{\bm{B}}_{1}^{\top}=-\nabla{\bm{A}}_{1}({\bm{R}}^{-\top}{\bm{P}}_{2}{\bm{R}}^{-1}){\bm{B}}_{1}^{\top}=\delta{\bm{A}}_{2}{\bm{B}}_{2}^{\top}$
$\implies-\nabla{\bm{A}}_{1}{\bm{P}}_{1}^{\top}{\bm{B}}_{1}^{\top}=-\nabla{\bm{A}}_{1}({\bm{R}}^{-\top}{\bm{P}}_{2}{\bm{R}}^{-1}){\bm{B}}_{1}^{\top}.$
@@
For any arbitrary ${\bm{R}}$, $\nabla{\bm{A}}_{1}$ and ${\bm{B}}_{1}$, there might not exist a pair of diagonal preconditioners ${\bm{P}}_{1},\ {\bm{P}}_{2}$ such that the above condition is satisfied, because ${\bm{R}}$ can be chosen in an adversarial fashion, such that ${\bm{R}}^{-\top}{\bm{P}}_{2}{\bm{R}}^{-1}$ is non-diagonal and distant from the left hand side. Consequently, we it is necessary to adopt matrix preconditioning to achieve transformation invariance.
@@ 3.2 Achieving Transformation Invariance
Note that the gradients of ${\bm{A}}$ and ${\bm{B}}$,
$\nabla{\bm{A}}=\nabla{\bm{Z}}{\bm{B}}\quad\text{and}\quad\nabla{\bm{B}}=\nabla{\bm{Z}}^{\top}{\bm{A}},$
depend on both the basis and the magnitude. To achieve transformation invariance, we introduce the concept of ”unmagnified gradients” distinguished from standard gradients by the symbol $\bar{\nabla}$:
¯$\bar{\nabla}{\bm{A}}\coloneqq\nabla{\bm{Z}}{\bm{U}}_{\bm{B}}=\nabla{\bm{A}}{\bm{R}}_{{\bm{B}}}^{-1},\quad\bar{\nabla}{\bm{B}}\coloneqq\nabla{\bm{Z}}^{\top}{\bm{U}}_{\bm{A}}=\nabla{\bm{B}}{\bm{R}}_{{\bm{A}}}^{-1}.$ (11)
These unmagnified gradients, that rely solely on the column spaces of ${\bm{A}}$ and ${\bm{B}}$, remain invariant to transformations of the LoRA factors. This invariance forms the cornerstone of our algorithm’s ability to achieve transformation invariance.
@@
¯$\bar{\nabla}{{\bm{A}}}(\bar{\nabla}{{\bm{A}}}^{\top}\bar{\nabla}{{\bm{A}}})^{-1/2}$
@@
The second part ${\bm{R}}_{{\bm{B}}}^{-\top}$ adjusts the magnitude of the update for different LoRA pairs. Since ${\bm{R}}_{{\bm{B}}}^{-\top}{\bm{B}}^{\top}={\bm{U}}_{{\bm{B}}}^{\top}$, this effectively takes out the magnitude of ${\bm{B}}^{\top}$ in $\delta{\bm{A}}{\bm{B}}^{\top}$, which is the first term in condition (4). We thus have
$\delta{\bm{A}}_{1}{\bm{B}}_{1}^{\top}$ =¯$=-\bar{\nabla}{\bm{A}}_{1}(\bar{\nabla}{{\bm{A}}_{1}}^{\top}\bar{\nabla}{{\bm{A}}_{1}})^{-1/2}({\bm{R}}_{{\bm{B}}_{1}}^{-\top}{\bm{B}}_{1}^{\top})=-\bar{\nabla}{\bm{A}}_{1}(\bar{\nabla}{{\bm{A}}_{1}}^{\top}\bar{\nabla}{{\bm{A}}_{1}})^{-1/2}{\bm{U}}_{{\bm{B}}_{1}}^{\top}$ (13)
=¯$=-\bar{\nabla}{\bm{A}}_{2}(\bar{\nabla}{{\bm{A}}_{2}}^{\top}\bar{\nabla}{{\bm{A}}_{2}})^{-1/2}{\bm{U}}_{{\bm{B}}_{2}}^{\top}=\delta{\bm{A}}_{2}{\bm{B}}_{2}^{\top}.$
Thus, LoRA-RITE update (17) satisfies condition (4) demonstrating that our proposed method satisfies transformation invariance. Note that this simplified update rule does not yet incorporate accumulated first and second moments, which will be addressed in the following paragraphs.
@@
indicating a potential loss of information from the accumulated second moment. To quantify this loss, for symmetric positive definite matrices ${\bm{X}}_{1},{\bm{X}}_{2}\in\mathbb{R}^{r\times r}$, we define
$d_{\lambda}({\bm{X}}_{1},{\bm{X}}_{2})\equiv\max_{i}|\lambda_{i}({\bm{X}}_{1})-\lambda_{i}({\bm{X}}_{2})|\leq\min_{\bm{U}}\|{\bm{X}}_{1}-{\bm{U}}{\bm{X}}_{2}{\bm{U}}^{\top}\|,$
where $\lambda_{i}({\bm{X}})$ is the $i$-th eigenvalue of ${\bm{X}}$, and ${\bm{U}}\in\mathbb{R}^{r\times r}$ is an orthogonal matrix that reflects our freedom to choose the basis. We then define the “escaped mass” as
@@
To compensate for this, we add $\rho_{{\bm{A}}_{t}}{\bm{I}}$ to our preconditioner, ensuring that
$\bar{{\bm{V}}}_{{\bm{A}}_{t}}+\rho_{{\bm{A}}_{t}}{\bm{I}}$
monotonically increases under a suitable choice of basis, even though the choice of basis does not influence the actual update.
@@
$\bar{{\bm{M}}}_{{\bm{A}}_{t}}=\beta_{1}\bar{{\bm{M}}}_{{\bm{A}}_{t-1}}{\bm{P}}_{{\bm{A}}_{t}}^{\top}+(1-\beta_{1})\bar{{\bm{S}}}_{{\bm{A}}_{t}}.$
@@
4:     Compute QR decomposition of the LoRA factor ${\bm{B}}_{t}$: ${\bm{B}}_{t}={\bm{U}}_{{\bm{B}}_{t}}{\bm{R}}_{{\bm{B}}_{t}}$;
5:     Compute the unmagnified gradient ¯$\bar{\nabla}{\bm{A}}_{t}=\nabla{\bm{A}}_{t}{\bm{R}}_{{\bm{B}}_{t}}^{-1}$ and ${\bm{P}}_{{\bm{A}}_{t}}=({\bm{U}}_{{\bm{B}}_{t}})^{\top}{\bm{U}}_{{\bm{B}}_{t-1}}$;
6:     Update the unmagnified second moment $\bar{{\bm{V}}}_{{\bm{A}}_{t}}={{\bm{P}}_{{\bm{A}}_{t}}}\bar{{\bm{V}}}_{{\bm{A}}_{t-1}}{\bm{P}}_{{\bm{A}}_{t}}^{\top}+(\bar{\nabla}{{\bm{A}}_{t}})^{\top}\bar{\nabla}{{\bm{A}}_{t}}$;
@@
10:     Update model parameters ${\bm{A}}_{t+1}={\bm{A}}_{t}-\eta_{t}\bar{{\bm{M}}}_{{\bm{A}}_{t}}{\bm{R}}_{{\bm{B}}}^{-\top}$.
@@
The time and space complexity of our algorithm is similar to first order methods like Adam when $r\ll m,n$. In each iteration of Algorithm 1, the dominant computational costs arise from (1) QR-decomposition for $m$-by-$r$ and $n$-by-$r$ matrices which takes $\mathcal{O}(nr^{2}+mr^{2})$ time, (2) matrix inverses and roots for $r$-by-$r$ matrices which takes $\mathcal{O}(r^{3})$ time, and (3) matmuls with time complexity $\mathcal{O}(nr^{2}+mr^{2})$. Thus, the overall complexity per step is $\mathcal{O}(mr^{2}+nr^{2}+r^{3})$. It is only $r$ times slower than Adam, and since $r$ is very small, this overhead is negligible when comparing with the back-propagating time. The memory cost of our method is $\mathcal{O}(mr+nr)$ which is the same as Adam. We summarize the time and space complexity of our method versus some commonly used optimizers in Table 7 in the Appendix.

### v2

@@ 3.1 Diagonal preconditioning is not enough for transformation invariance
In this section we show that diagonal preconditioning is not enough to achieve transformation invariance. Recall that the LoRA factors are ${\bm{A}}\in\mathbb{R}^{m\times r}$ and ${\bm{B}}\in\mathbb{R}^{n\times r}$. When updating ${\bm{A}}$, most existing optimizers utilize the following update:
$\operatorname{vec}(\delta{\bm{A}})=-\eta{\bm{X}}\operatorname{vec}(\nabla{\bm{A}}),$ (7)
@@
where $\operatorname{vec}(\cdot)$ lists the elements of a matrix as a vector in column-major order and ${\bm{X}}\in\mathbb{R}^{mr\times mr}$ is a symmetric positive definite preconditioning matrix. Diagonal preconditioning methods like Adam and Adagrad assume ${\bm{X}}$ is a diagonal matrix, while matrix preconditioning methods such as Shampoo and CASPR use non-diagonal ${\bm{X}}$.
@@
Consider the LoRA pairs $({\bm{A}}_{1},{\bm{B}}_{1})$ and $({\bm{A}}_{2},{\bm{B}}_{2})$ such that: ${\bm{A}}_{2}={\bm{A}}_{1}{\bm{R}},~{}{\bm{B}}_{2}={\bm{B}}_{1}{\bm{R}}^{-\top}$, where ${\bm{R}}$ is an invertible matrix. By (4), the gradient with respect to the LoRA factors are:
$\nabla{\bm{A}}_{2}=\nabla{\bm{Z}}{\bm{B}}_{2}=\nabla{\bm{A}}_{1}{\bm{R}}^{-\top},\quad\nabla{\bm{B}}_{2}=\nabla{\bm{Z}}^{\top}{\bm{A}}_{2}=\nabla{\bm{B}}_{1}{\bm{R}}.$ (8)
@@
Let us consider the case when $n=m=1$. For ${\bm{A}}\in\mathbb{R}^{1\times r}$, we can substitute $\operatorname{vec}(\delta{\bm{A}})=\delta{\bm{A}}^{\top}$ and $\operatorname{vec}(\nabla{\bm{A}})=\nabla{\bm{A}}^{\top}$ in update (7) and apply transpose on both sides to get:
$\delta{\bm{A}}=-\eta\nabla{\bm{A}}{\bm{X}}^{\top}=-\eta\nabla{\bm{Z}}{\bm{B}}{\bm{X}},\quad\delta{\bm{B}}=-\eta\nabla{\bm{B}}{\bm{X}}^{\top}=-\eta\nabla{\bm{Z}}^{\top}{\bm{A}}{\bm{X}}^{\top}$ (9)
@@
In this case, if we have two equivalent LoRA pairs $({\bm{A}}_{1},{\bm{B}}_{1}),({\bm{A}}_{2},{\bm{B}}_{2})$ with their corresponding preconditioners ${\bm{X}}_{1},{\bm{X}}_{2}$, then from (8) and (9), we have
$\delta{\bm{A}}_{1}{\bm{B}}_{1}^{\top}+{\bm{A}}_{1}\delta{\bm{B}}_{1}^{\top}+\delta{\bm{A}}_{1}\delta{\bm{B}}_{1}^{\top}=-\eta\nabla{\bm{Z}}{\bm{B}}_{1}{\bm{X}}_{1}^{\top}{\bm{B}}_{1}^{\top}-\eta{\bm{A}}_{1}{\bm{X}}_{1}{\bm{A}}_{1}^{\top}\nabla{\bm{Z}}+\eta^{2}\nabla{\bm{Z}}{\bm{B}}_{1}{\bm{X}}_{1}^{\top}{\bm{X}}_{1}{\bm{A}}_{1}^{\top}\nabla{\bm{Z}}$ (10)
and
$\delta{\bm{A}}_{2}{\bm{B}}_{2}^{\top}+{\bm{A}}_{2}\delta{\bm{B}}_{2}^{\top}+\delta{\bm{A}}_{2}\delta{\bm{B}}_{2}^{\top}=-\eta\nabla{\bm{Z}}{\bm{B}}_{2}{\bm{X}}_{2}^{\top}{\bm{B}}_{2}^{\top}-\eta{\bm{A}}_{2}{\bm{X}}_{2}{\bm{A}}_{2}^{\top}\nabla{\bm{Z}}+\eta^{2}\nabla{\bm{Z}}{\bm{B}}_{2}{\bm{X}}_{2}^{\top}{\bm{X}}_{2}{\bm{A}}_{2}^{\top}\nabla{\bm{Z}}$ (11)
$=$ $-\eta\nabla{\bm{Z}}{\bm{B}}_{1}({\bm{R}}^{-\top}{\bm{X}}_{2}^{\top}{\bm{R}}^{-1}){\bm{B}}_{1}^{\top}-\eta{\bm{A}}_{1}({\bm{R}}{\bm{X}}_{2}{\bm{R}}^{\top}){\bm{A}}_{1}^{\top}\nabla{\bm{Z}}+\eta^{2}\nabla{\bm{Z}}{\bm{B}}_{1}{\bm{R}}^{-\top}{\bm{X}}_{2}^{\top}{\bm{X}}_{2}{\bm{R}}^{\top}{\bm{A}}_{1}^{\top}\nabla{\bm{Z}}.$
For (10) and (11) to be equal for any arbitrary $\eta$, one can see that it is necessary to have
$\nabla{\bm{Z}}{\bm{B}}_{1}{\bm{X}}_{1}^{\top}{\bm{X}}_{1}{\bm{A}}_{1}^{\top}\nabla{\bm{Z}}=\nabla{\bm{Z}}{\bm{B}}_{1}{\bm{R}}^{-\top}{\bm{X}}_{2}^{\top}{\bm{X}}_{2}{\bm{R}}^{\top}{\bm{A}}_{1}^{\top}\nabla{\bm{Z}},$ (12)
@@
as it is the only term quadratic to $\eta$. However, there might not exist a pair of diagonal preconditioners ${\bm{X}}_{1},\ {\bm{X}}_{2}\neq\bm{0}$ such that the above condition is satisfied for arbitrary $\nabla{\bm{Z}},{\bm{A}},$ and ${\bm{B}}$, because ${\bm{R}}$ can be a full matrix chosen in an adversarial fashion, such that ${\bm{R}}^{-\top}{\bm{X}}_{2}^{\top}{\bm{X}}_{2}{\bm{R}}^{\top}$ is non-diagonal and thus different from the left hand side. Consequently, we conclude it is necessary to adopt matrix preconditioning to achieve transformation invariance.
@@ 3.2 Achieving Transformation Invariance
Note that the gradients of ${\bm{A}}$ and ${\bm{B}}$, $\nabla{\bm{A}}=\nabla{\bm{Z}}{\bm{B}}~{}\text{and}~{}\nabla{\bm{B}}=\nabla{\bm{Z}}^{\top}{\bm{A}}$, depend on both the basis and the magnitude. To achieve transformation invariance, we introduce the concept of “unmagnified gradients” distinguished from standard gradients by the symbol $\bar{\nabla}$:
¯$\bar{\nabla}{\bm{A}}\coloneqq\nabla{\bm{Z}}{\bm{U}}_{\bm{B}}=\nabla{\bm{A}}{\bm{R}}_{{\bm{B}}}^{\dagger},\quad\bar{\nabla}{\bm{B}}\coloneqq\nabla{\bm{Z}}^{\top}{\bm{U}}_{\bm{A}}=\nabla{\bm{B}}{\bm{R}}_{{\bm{A}}}^{\dagger},$ (13)
where ${\bm{R}}_{{\bm{A}}}^{\dagger}$ and ${\bm{R}}_{{\bm{B}}}^{\dagger}$ are the pseudo-inverse of ${\bm{R}}_{{\bm{A}}}$ and ${\bm{R}}_{{\bm{B}}}$ respectively. These unmagnified gradients, relying solely on the column spaces of ${\bm{A}}$ and ${\bm{B}}$, remain invariant to transformations of the LoRA factors. This invariance forms the cornerstone of our algorithm’s ability to achieve transformation invariance.
@@
$-\eta\bar{\nabla}{{\bm{A}}}(\bar{\nabla}{{\bm{A}}}^{\top}\bar{\nabla}{{\bm{A}}})^{-1/2}$
@@
The second part $({\bm{R}}_{{\bm{B}}}^{\top})^{\dagger}$ adjusts the magnitude of the update for different LoRA pairs. Since $({\bm{R}}_{{\bm{B}}}^{\top})^{\dagger}{\bm{B}}^{\top}={\bm{U}}_{{\bm{B}}}^{\top}$, this effectively takes out the magnitude of ${\bm{B}}^{\top}$ in $\delta{\bm{A}}{\bm{B}}^{\top}$. We thus have
$\delta{\bm{A}}_{2}$ $=-\eta\bar{\nabla}{\bm{A}}_{2}(\bar{\nabla}{{\bm{A}}_{2}}^{\top}\bar{\nabla}{{\bm{A}}_{2}})^{-1/2}({\bm{R}}_{{\bm{B}}_{2}}^{\top})^{\dagger}=-\eta\bar{\nabla}{\bm{A}}_{2}(\bar{\nabla}{{\bm{A}}_{2}}^{\top}\bar{\nabla}{{\bm{A}}_{2}})^{-1/2}{\bm{U}}_{{\bm{B}}_{2}}^{\top}({\bm{B}}_{2}^{\top})^{\dagger}$ (15)
$=-\eta{\nabla}{\bm{Z}}{\bm{U}}_{{\bm{B}}_{2}}(\bar{\nabla}{{\bm{A}}_{2}}^{\top}\bar{\nabla}{{\bm{A}}_{2}})^{-1/2}{\bm{U}}_{{\bm{B}}_{2}}^{\top}({\bm{B}}_{2}^{\top})^{\dagger}=-\eta{\nabla}{\bm{Z}}({\bm{U}}_{{\bm{B}}_{2}}\bar{\nabla}{{\bm{A}}_{2}}^{\top}\bar{\nabla}{{\bm{A}}_{2}}{\bm{U}}_{{\bm{B}}_{2}}^{\top})^{-1/2}({\bm{B}}_{2}^{\top})^{\dagger}$
$=-\eta{\nabla}{\bm{Z}}({\bm{U}}_{{\bm{B}}_{1}}\bar{\nabla}{{\bm{A}}_{1}}^{\top}\bar{\nabla}{{\bm{A}}_{1}}{\bm{U}}_{{\bm{B}}_{1}}^{\top})^{-1/2}({\bm{B}}_{2}^{\top})^{\dagger}=\delta{\bm{A}}_{1}{\bm{B}}_{1}^{\top}({\bm{B}}_{2}^{\top})^{\dagger}.$
Similarly, $\delta{\bm{B}}_{2}=\delta{\bm{B}}_{1}{\bm{A}}_{1}^{\top}({\bm{A}}_{2}^{\top})^{\dagger}$. Consequently, we have
${\bm{A}}_{1}{\bm{B}}_{1}^{\top}+\delta{\bm{A}}_{1}{\bm{B}}_{1}^{\top}+{\bm{A}}_{1}\delta{\bm{B}}_{1}^{\top}+\delta{\bm{A}}_{1}\delta{\bm{B}}_{1}^{\top}={\bm{A}}_{2}{\bm{B}}_{2}^{\top}+\delta{\bm{A}}_{2}{\bm{B}}_{2}^{\top}+{\bm{A}}_{2}\delta{\bm{B}}_{2}^{\top}+\delta{\bm{A}}_{2}\delta{\bm{B}}_{2}^{\top}.$
This demonstrates that our proposed method satisfies transformation invariance. Note that this simplified update rule does not yet incorporate accumulated first and second moments, which will be addressed in the following paragraphs.
@@
indicating a potential loss of information from the accumulated second moment. To quantify this loss, for symmetric positive definite matrices ${\bm{E}}_{1},{\bm{E}}_{2}\in\mathbb{R}^{r\times r}$, we define
$d_{\lambda}({\bm{E}}_{1},{\bm{E}}_{2})\equiv\max_{i}|\lambda_{i}({\bm{E}}_{1})-\lambda_{i}({\bm{E}}_{2})|\leq\min_{\bm{U}}\|{\bm{E}}_{1}-{\bm{U}}{\bm{E}}_{2}{\bm{U}}^{\top}\|,$
where $\lambda_{i}({\bm{E}})$ is the $i$-th eigenvalue of ${\bm{E}}$, and ${\bm{U}}\in\mathbb{R}^{r\times r}$ is an orthogonal matrix that reflects our freedom to choose the basis. We then define the “escaped mass” as
@@
To compensate for this, we add $\rho_{{\bm{A}}_{t}}{\bm{I}}$ to our preconditioner, ensuring that $\bar{{\bm{V}}}_{{\bm{A}}_{t}}+\rho_{{\bm{A}}_{t}}{\bm{I}}$ monotonically increases under a suitable choice of basis, even though the choice of basis does not influence the actual update.
@@
$\hskip 30.0pt\quad\quad\bar{{\bm{M}}}_{{\bm{A}}_{t}}=\beta_{1}\bar{{\bm{M}}}_{{\bm{A}}_{t-1}}{\bm{P}}_{{\bm{A}}_{t}}^{\top}+(1-\beta_{1})\bar{{\bm{S}}}_{{\bm{A}}_{t}}.$
@@
4:     Compute polar decomposition of the LoRA factor ${\bm{B}}_{t}$: ${\bm{B}}_{t}={\bm{U}}_{{\bm{B}}_{t}}{\bm{R}}_{{\bm{B}}_{t}}$;
5:     Compute the unmagnified gradient ¯$\bar{\nabla}{\bm{A}}_{t}=\nabla{\bm{A}}_{t}{\bm{R}}_{{\bm{B}}_{t}}^{\dagger}$ and ${\bm{P}}_{{\bm{A}}_{t}}=({\bm{U}}_{{\bm{B}}_{t}})^{\top}{\bm{U}}_{{\bm{B}}_{t-1}}$;
6:     Update the unmagnified second moment $\bar{{\bm{V}}}_{{\bm{A}}_{t}}={{\bm{P}}_{{\bm{A}}_{t}}}\bar{{\bm{V}}}_{{\bm{A}}_{t-1}}{\bm{P}}_{{\bm{A}}_{t}}^{\top}+(\bar{\nabla}{{\bm{A}}_{t}})^{\top}\bar{\nabla}{{\bm{A}}_{t}}/m$;
@@
10:     Update model parameters ${\bm{A}}_{t+1}={\bm{A}}_{t}-\eta_{t}\bar{{\bm{M}}}_{{\bm{A}}_{t}}({\bm{R}}_{{\bm{B}}}^{\top})^{\dagger}$.
@@
The time and space complexity of our algorithm is similar to first order methods like Adam when $r\ll m,n$. In each iteration of Algorithm 1, the dominant computational costs arise from (1) polar-decomposition for $m$-by-$r$ and $n$-by-$r$ matrices which takes $\mathcal{O}(nr^{2}+mr^{2})$ time, (2) matrix inverses and roots for $r$-by-$r$ matrices which takes $\mathcal{O}(r^{3})$ time, and (3) matmuls with time complexity $\mathcal{O}(nr^{2}+mr^{2})$. Thus, the overall complexity per step is $\mathcal{O}(mr^{2}+nr^{2}+r^{3})$. It is only $r$ times slower than Adam, and since $r$ is very small, this overhead is negligible when comparing with the back-propagating time. The memory cost of our method is $\mathcal{O}(mr+nr)$ which is the same as Adam. We summarize the time and space complexity of our method versus some commonly used optimizers in Table 7 in the Appendix.